You have to slove the system of linear equations using linear combinations.
5x-3y=26
3x+3y=8

Answer by stanbon(48551)

5x-3y=26
3x+3y=8
------------
Add the equations to get:
8x=34
x=17/4
Substitute into 3x+3y=8 to solve for "y", as follows:
3(17/4)+3y=8
51+12y=32
12y=15
y=5/4
Solution: x=17/4 y=5/4
Cheers,
Stan H.